OpenStudy (anonymous):

An Olympic marathoner passes the 5km point of the race after 15 minutes and reaches the 22km point after 3.1hours. find the average speed of the runner in km/hr. help please

OpenStudy (anonymous):

avr=[(5*1/4)+(22*3.1)]/(3.1+1/4)=20.73km/hr

OpenStudy (anonymous):

it says thats wrong...

OpenStudy (anonymous):

8.06km/ hr

OpenStudy (anonymous):

the average speed = total distance / time if we are talking about the average for the whole 22 km it is 22 / 3.1 = 7.1 km/hr for the second part, distance = 22 - 5 = 17 and time = 3.1 - .25 = 2.85 avg speed = 17/ 2.85 = 5.96